OpenStudy (anonymous):

I need help, First question ever asked here, hope this helps! Solve for y, then find the value of y when given x=-2 6x=7-4y I tried going at it from all ends. How do I work through it?

OpenStudy (anonymous):

When x = -2 6x = -12 So 12 = 7 - 4y hence ... can you take it from here?

OpenStudy (johnweldon1993):

Solve for 'y first I suppose \[\large 6x = 7 - 4y\] Subtract 7 from both sides of the equation \[\large 6x - 7 = -4y\] Divide both sides by -4 \[\large y = \frac{6x - 7}{-4}\] Now just plug in x = -2 \[\large y = \frac{6(-2) - 7}{-4} = \frac{-12 - 7}{-4} = ?\]

OpenStudy (anonymous):

Thanks Guys, helped a lot!